OpenStudy (anonymous):

For what values of x does x2 +10x + 2 = 0

OpenStudy (p0sitr0n):

\[x1,2 =(-b \pm \sqrt{b^2-4ac})/2a\], where a is the term before x^2 (1), b= 10, c =2

OpenStudy (p0sitr0n):

\[x1,2 = -5\pm \sqrt{92}/2\]